### v1

@@ 2.6 A Tree Cut-Sparsifier of Quality $O(\log^{2}n\log\log n)$
Following the naming of [harrelson2003polynomial, RackeS14], we call the partitioning algorithm from Section 2.5, that takes $S$ and returns $\{S_{i}\}_{i=1}^{t}$, the merge phase. As explained in Section 2.5, this phase does not guarantee that the inter-cluster edges $F$ and the boundary edges $B$ are expanding together, just that the edges of $F$ are expanding, and we can route from $F$ to $B$. Recall from Section 2.4 that because the distribution of demands on the boundary edges is not uniform, source and target demands do not cancel-out, so the load on the boundary edges accumulates as we go up along the tree. This degrades the quality of the tree cut-sparsifier by an $O(\log n)$ factor. To get the improved Theorem 2.8, in Section 6 we show how to implement a version of the refinement phase (see Section 2.4) in near-linear time. Specifically, we show that given a cluster $S$ we can partition it to sub-clusters $\{S_{1},\ldots,S_{t}\}$ such that each sub-cluster respects the all-to-all multi-commodity flow problem on its boundary edges. This is a slightly weaker requirement than the precondition of [bienkowski2003practical] and the requirement of [harrelson2003polynomial, RackeS14]. Additionally, we show that we can route from the inter-cluster edges of this refinement partition to the boundary of $S$. To perform this partition in near-linear time, we again use the result of [ADK25] which allows us to either find a balanced sparse cut (with respect to the all-to-all flow problem on the boundary of the current sub-cluster) or ensure that the boundary edges are expanding. Additionally, it lets us route from the sparse cut’s edges to the boundary edges of one of the sides of the cut. Crucially, however, contrary to the merge phase, we get no guarantee that the size $|S_{i}|$ is smaller than $|S|$ by a constant factor (and in fact it may be that the partition is simply $\{S\}$).
@@
Using the refinement phase, we amend the hierarchical decomposition tree by performing this refinement after each call to the merge phase algorithm. This gives a decomposition which alternates between a refinement phase partition, that allows to spread the demand uniformly on the boundary, and a merge phase partition that shrinks the clusters. Because each sub-cluster $S_{i}$ in the refinement phase respects the all-to-all demand matrix on its boundary edges, we can move the demands to a uniform distribution on these boundary edges. Furthermore, since $S_{i}\in V_{T}$ (and by assumption $Q$ is routable in $T$), the cut $E(S_{i},V\setminus S_{i})$ respects $Q$, we get that at most $1$ unit of mass per edge needs to leave the cluster. This allows us to “reset” the load on each boundary edge to be $1$ and avoid accumulating the loads on the boundary edges as we go up the tree. By setting $\uptau=1$ in the merge phase, we improve the degradation (in quality) incurred by the merge phase by a factor of $O(\log n)$, to be $O(\log^{2}n)$ (in total over all sub-clusters), instead of $O(\log^{3}n)$. This is because we pay $O(\frac{\log n}{\uptau})=O(\log n)$, plus an additional $O(\log n)$ term for the cut-matching component like before. As for the degradation in quality incurred by the refinement phase, we adapt the analysis of [harrelson2003polynomial, RackeS14] to show that it is bounded by $O(\log^{2}n\log\log n)$ (in total over all sub-clusters), which gives Theorem 2.8. In particular, our tree is also a tree flow-sparsifier of quality $O(\log^{3}n\log\log n)$, which gives Theorem 2.11.
@@ 3.2 Expansion
So far, expansion has been defined as a property of a graph. In many of our later arguments, however, we talk about individual subsets $A\subseteq V$ that expand well within a given graph $G$ under some weighting. The next definition captures this idea: it says that a subset AA $\alpha$-expands with respect to $\mu$ if no cut of $V$ separates $A$ too unevenly, when the imbalance is measured according to the $\mu$-measure of the vertices. This notion generalizes over the $\alpha$-cut-linked notion of [chekuri2005multicommodity].
@@
In other words, if we let $\mu^{\prime}=\mu\cdot\mu_{A}$, then $A\subseteq V$, $\alpha$-expands in $G$ with respect to $\mu$, if and only if $G$ is an $\alpha$-expander with respect to $\mu^{\prime}$.
When $\mu\equiv 1$, we just say that AA $\alpha$-expands in $G$.
@@
The next remark bridges the two viewpoints introduced in Definitions 2.6 and 3.5: the notion of a set $A$ that $\alpha$-expands with respect to a measure $\mu$, and the notion of a graph $G$ that $\alpha$-respects a corresponding flow problem defined by $\mu$.
@@
In particular (by setting $\mu\equiv\mathbbm{1}$), up to constant factors, AA $\alpha$-expands in $G$ is equivalent to $G$ $\alpha$-respecting the all-to-all flow problem on $A$ (recall Definition 2.2).
@@ 3.3 Cut-Matching Game
The following is a recent result of [ADK25], which generalizes [agassy2023expander]. They show that in near-linear time we can get a balanced sparse cut, or certify that a large part of the graph is an expander, with respect to a general vertex measure. [^4] Theorem 3.9 requires that $\mu(v)$ will be in $\left[\frac{1}{\operatorname{poly}(n)},\operatorname{poly}(n)\right]\cup\{0\}$, which will hold for our usages.
Theorem 3.9 ([ADK25, Theorem 5.2]).
@@
The authors of [ADK25] supply additional guarantees in Cases (2) and (3) of Theorem 3.9 as follows. First, in both cases $\mu(A)\geq\frac{1}{4}\mu(V)$.
Additionally, in Case (2), there exist constant $c>0$ and a sequence of cuts $S_{t}\subseteq A_{t}$, $0\leq t<T^{\prime}$ with $T^{\prime}=O(\log^{2}n)$, such that $A_{0}=V$, $A_{t+1}=A_{t}\setminus S_{t}$ and $\mu(\bigcup_{t=0}^{T^{\prime}-1}{S_{t}})=O(\frac{\mu(V)}{\log n})$. Each $S_{t}$ is sparse: $\Phi^{\mu}_{G[A_{t}]}(S_{t},A_{t+1})\leq c\phi\log n$. The returned cut is simply $\bar{A}=\bigcup_{t}S_{t}$. Moreover, [ADK25, Corollary A.3] gives
  • There exists a flow routable with congestion $O(1)$ in $G[S_{t}]$, such that each cut edge of $E(S_{t},A_{t}\setminus S_{t})$ is a source of one unit of flow and each vertex of $S_{t}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
  • There exists a flow routable with congestion $O(1)$ in $G[A_{t}\setminus S_{t}]$, such that each cut edge of $E(S_{t},A_{t}\setminus S_{t})$ is a source of one unit of flow and each vertex of $A_{t}\setminus S_{t}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
In other words, in each step we can route flow from the cut edges $E(S_{t},A_{t}\setminus S_{t})$ to either the side $S_{t}$ or $A_{t}\setminus S_{t}$.
In Case (3), this sequence of cuts exist as well, but they run a “trimming step” to get some larger cut $\bar{A}\supseteq\bigcup_{t}S_{t}$ which is returned. Additionally, [ADK25, Corollary A.5] gives
  • There exists a flow routable with congestion $O(1)$ in $G[A]$, such that each cut edge of $E(A,\bar{A})$ is a source of one unit of flow and each vertex of $A$ receives at most $O(\phi\cdot\mu(v))$ units.
  • There exists a flow routable with congestion $O(1)$ in $G[\bar{A}]$, such that each cut edge of $E(A,\bar{A})$ is a source of one unit of flow and each vertex of $\bar{A}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
When we use Theorem 3.9, when Case (3) occurs, we sometimes prefer to have the guarantee that $\mu(\bar{A})$ is significantly smaller than $\mu(A)$. This can be done by simply terminating with Case (2) if $\mu(\bar{A})$ and $\mu(A)$ are balanced. This is captured by the following Corollary, which is a straightforward strengthening of Theorem 3.9. The only difference is the bound on $\mu(\bar{A})$ in Case (3).
Corollary 3.11 ([ADK25, Corollary 5.3]).
  1. 1.
  2. 2.
  3. 3.
    We find a cut $(A,\bar{A})$ with $0<\mu(\bar{A})\leq\frac{\mu(V)}{\log n}$, $\Phi^{\mu}_{G}(A,\bar{A})=O(\phi\log n)$, and with high probability, $G[A]$ is an $\Omega(\phi)$-expander with respect to $\mu$.
Proof.
If Theorem 3.9 terminated with Case (3) but $\frac{\mu(V)}{\log n}<\mu(\bar{A})\leq\frac{\mu(V)}{2}$, we simply terminate with Case (2) instead. ∎
We emphasize that in Corollary 3.11, some of the flows given by Remark 3.10 in Case (2) may no longer exist. Specifically, in Case (2) the cut is no longer given directly as $\bar{A}=\bigcup_{t}{S_{t}}$, because the trimming step was applied on $\bigcup_{t}{S_{t}}$.
Remark 3.12.
In both Theorem 3.9 and Corollary 3.11, in Case (3), the fact that $G[A]$ is an $\Omega(\phi)$-expander with respect to $\mu$, implies that AA $\Omega(\phi)$-expands with respect to $\mu$ in $G$ by Fact 3.7.
@@ 3.4 Demand States
When we describe the solution to a multi-commodity flow problem, it will be useful to talk about intermediate stages of the flow. Analogously, we wish to talk about intermediate states of the distribution of demands when showing that a graph respects a given demand matrix. Intuitively, we imagine the vertices of the graph holding some mass (demand) of each commodity, which can be positive or negative. These demands move from one vertex to another according to demand matrices (i.e., multi-commodity flow problems), incurring some congestion (or, in the cut-sparsifier case, charge, as detailed in Section 5) on the edges of the graph. When a positive demand meets a negative demand of the same commodity, they cancel out. This scheme is made formal in the definitions in this section.
@@
Consider a set of commodities $\mathcal{K}$. We consider for each vertex $v\in V$ a demand state on $v$, which is a vector $P(v)\in\mathop{\mathbb{R}}^{\mathcal{K}}$, where $P(v,k)$, which may be either positive or negative, defines the amount of commodity $k\in\mathcal{K}$ which resides at $v\in V$ and needs to be shipped out. We denote by $\left\|P(v)\right\|_{1}=\sum_{k\in\mathcal{K}}{|P(v,k)|}$ the load on $v$.
A demand state on $G$ (or simply a demand state when $G$ is clear from context) is a collection of demand states on each $v\in V$, i.e., $P\in\mathop{\mathbb{R}}^{V\times\mathcal{K}}$, where $P(v,\cdot)\mathrel{\mathop{:}}=P(v)$ is $v$’s demand state.
@@
$\operatorname{dem}_{P}(A,B)\mathrel{\mathop{:}}=\sum_{k\in\mathcal{K}}{|\sum_{u\in A}{P(u,k)}|}=\sum_{k\in\mathcal{K}}{|\sum_{u\in B}{P(u,k)}|}.$
@@
Note that the notion of a valid demand state generalizes the notion of a demand matrix from Definition 2.1. Indeed, a demand matrix is a $V\times V$ matrix $Q$ where $Q(u,v)$ is the amount of $u$’s commodity that has to go to $v$. Such a specification corresponds to a valid demand state $P$ in which $\mathcal{K}=V$, and the amount shipped out of $u$ of $u$’s commodity is $P(u,u)=\sum_{v\in V\setminus\{u\}}{Q(u,v)}$, and the amount shipped to $v$ of $u$’s commodity is P(v,u)=Q(u,v)P(v,u)=-Q(u,v), $v\neq u$. The standard definitions of congestion and demand across a cut of $Q$ equal those of $P$.
On the other hand, given a valid demand state, we can define the corresponding multi-commodity flow problem (according to Definition 2.1) by adding to the graph a source vertex and sink vertex corresponding to each commodity, with appropriate edge capacities according to the demands in $P$. This reduction will not be used in this paper, so we omit the full details.
@@
The following fact, which we use extensively, is a direct consequence of Definitions 3.14 and 3.16:
Fact 3.17.
@@
Suppose we have a demand state $P$. Given a demand matrix $Q$ (recall Definition 2.1 of a multi-commodity flow problem), we can “move” the demand vectors in $P$ according to the requirements of $Q$. Each vertex $u\in V$ will send $Q(u,v)$ units in total to $v\in V$, and we split the demands evenly according to the distribution of commodities in the demand vector $P(u)$. In other words, $\frac{P(u,k)}{\left\|P(u)\right\|_{1}}\cdot Q(u,v)$ units of commodity $k\in\mathcal{K}$ will be shipped to $v$. If $P(u,k)<0$ this corresponds to shipping units of commodity $k$ from $v$ to $u$. This results in a demand state $P^{\uparrow Q}$, called the updated demand state:
@@
Example. Suppose the demand state of a vertex $u$ is P(u)=(20,70,10)P(u)=(20,-70,10), representing three commodities. If a flow problem $Q$ routes $30$ units from $u$ to a vertex $v$, then the updated demand state moves $30\times(0.2,-0.7,0.1)$ units of demand from $u$ to $v$, leaving $u$ with a demand state of $P^{\uparrow Q}(u)=(14,-49,7)$.
We will make frequent use of the following fact. It states that because we only move the mass around, no mass of any commodity is created or destroyed. This means that the total mass of each commodity is the same in $P$ and $P^{\uparrow Q}$, so it is zero in $P-P^{\uparrow Q}$. Therefore, $P-P^{\uparrow Q}$ is a valid demand state. The full proof of the fact is deferred to Appendix A.
@@
Remark 3.20.
Let $P$ be a (not necessarily valid) demand state on the set of commodities $\mathcal{K}$. Of special interest are the following cases:
@@
    $Q$ is a scaled all-to-all flow on some subset $A\subseteq V$: $D_{Q}(u,v)=\frac{\left\|P(u)\right\|_{1}}{|A|}$ if $u,v\in A$, and otherwise $D_{Q}(u,v)=0$. I.e., instead of sending $1$ unit out of $u$ in total, like in the all-to-all case (see Definition 2.2), we send $\left\|P(u)\right\|_{1}$ units. In this case, because the demand vector $P(u)$ for $u\in A$, was spread uniformly among all vertices of $A$, the new demand vector of each vertex, $P^{\uparrow Q}(v)$ is the same for all $v\in A$ (i.e., it does not depend on $v$). Specifically, $P^{\uparrow Q}(v)=\frac{1}{|A|}\sum_{u\in A}{P(u)}$ for $v\in A$. This corresponds to the fact that mass was distributed uniformly.
@@
    $Q$ moves demands from a set $A\subseteq V$ to $V\setminus A$: For any $u\in A$, $D_{Q}(v,u)=0$ for all $v\in V$, and also $\sum_{v\in V\setminus A}{D_{Q}(u,v)}=\left\|P(u)\right\|_{1}$. In this case, note that $P^{\uparrow Q}$ is supported on $V\setminus A$.

### v2

@@ 2.6 A Tree Cut-Sparsifier of Quality $O(\log^{2}n\log\log n)$
Following the naming of [15, 28], we call the partitioning algorithm from Section 2.5, that takes $S$ and returns $\{S_{i}\}_{i=1}^{t}$, the merge phase. As explained in Section 2.5, this phase does not guarantee that the inter-cluster edges $F$ and the boundary edges $B$ are expanding together, just that the edges of $F$ are expanding, and we can route from $F$ to $B$. Recall from Section 2.4 that because the distribution of demands on the boundary edges is not uniform, source and target demands do not cancel-out, so the load on the boundary edges accumulates as we go up along the tree. This degrades the quality of the tree cut-sparsifier by an $O(\log n)$ factor. To get the improved Theorem 2.8, in Section 6 we show how to implement a version of the refinement phase (see Section 2.4) in near-linear time. Specifically, we show that given a cluster $S$ we can partition it to sub-clusters $\{S_{1},\ldots,S_{t}\}$ such that each sub-cluster respects the all-to-all flow problem on its boundary edges. This is a slightly weaker requirement than the precondition of [7] and the requirement of [15, 28]. Additionally, we show that we can route from the inter-cluster edges of this refinement partition to the boundary of $S$. To perform this partition in near-linear time, we again use the result of [2] which allows us to either find a balanced sparse cut (with respect to the all-to-all flow problem on the boundary of the current sub-cluster) or ensure that the boundary edges are expanding. Additionally, it lets us route from the sparse cut’s edges to the boundary edges of one of the sides of the cut. Crucially, however, contrary to the merge phase, we get no guarantee that the size $|S_{i}|$ is smaller than $|S|$ by a constant factor (and in fact it may be that the partition is simply $\{S\}$).
@@
Using the refinement phase, we amend the hierarchical decomposition tree by performing this refinement after each call to the merge phase algorithm. This gives a decomposition which alternates between a refinement phase partition, that allows to spread the demand uniformly on the boundary, and a merge phase partition that shrinks the clusters. Because each sub-cluster $S_{i}$ in the refinement phase respects the all-to-all demand matrix on its boundary edges, we can move the demands to a uniform distribution on these boundary edges. Furthermore, since $S_{i}\in V_{T}$ (and by assumption $Q$ is routable in $T$), the cut $E(S_{i},V\setminus S_{i})$ respects $Q$, we get that at most $1$ unit of mass per edge needs to leave the cluster. This allows us to “reset” the load on each boundary edge to be $1$ and avoid accumulating the loads on the boundary edges as we go up the tree. By setting $\uptau=1$ in the merge phase, we improve the degradation (in quality) incurred by the merge phase by a factor of $O(\log n)$, to be $O(\log^{2}n)$ (in total over all sub-clusters), instead of $O(\log^{3}n)$. This is because we pay $O(\frac{\log n}{\uptau})=O(\log n)$, plus an additional $O(\log n)$ term for the cut-matching component like before. As for the degradation in quality incurred by the refinement phase, we adapt the analysis of [15, 28] to show that it is bounded by $O(\log^{2}n\log\log n)$ (in total over all sub-clusters), which gives Theorem 2.8. In particular, our tree is also a tree flow-sparsifier of quality $O(\log^{3}n\log\log n)$, which gives Theorem 2.12.
@@ 3.2 Expansion
So far, expansion has been defined as a property of a graph. In many of our later arguments, however, we talk about individual subsets $A\subseteq V$ that expand well within a given graph $G$ under some weighting. The next definition captures this idea: it says that a subset $A$, $\alpha$-expands with respect to $\mu$, if no cut of $V$ separates $A$ too unevenly, when the imbalance is measured according to the $\mu$-measure of the vertices. This notion generalizes the $\alpha$-cut-linked notion of [10].
@@
In other words, if we let $\mu^{\prime}=\mu\cdot\mu_{A}$, then $A\subseteq V$, $\alpha$-expands in $G$ with respect to $\mu$, if and only if $G$ is an $\alpha$-expander with respect to $\mu^{\prime}$. When $\mu\equiv 1$, we just say that AA $\alpha$-expands in $G$.
@@
The next remark bridges the two viewpoints introduced in Definitions 2.6 and 3.5: the notion of a set $A$ that $\alpha$-expands with respect to a measure $\mu$, and the notion of a graph $G$ that $\alpha$-respects a corresponding demand matrix defined by $\mu$.
@@
In particular (by setting $\mu\equiv\mathbbm{1}$), if AA $\alpha$-expands in $G$, then $G$ $\frac{\alpha}{2}$-respects the all-to-all flow problem on $A$ (recall Definition 2.2), and conversely if $G$ $\alpha$-respects the all-to-all flow problem on $A$ then AA $\alpha$-expands in $G$.
@@ 3.3 Cut-Matching Game
The following is a recent result of [2], which generalizes [1]. They show that in near-linear time we can get a balanced sparse cut, or certify that a large part of the graph is an expander, with respect to a general vertex measure. [^5] Theorem 3.9 requires that $\mu(v)$ will be in $\left[\frac{1}{\operatorname{poly}(n)},\operatorname{poly}(n)\right]\cup\{0\}$, which will hold for our usages.
Convention (orientation of cuts).
In all balanced-cut results we use (Theorems 3.9 and 3.10, and Corollary 6.5), we orient a cut $(A,\bar{A})$ so that the left side $A$ is conceptually the larger side with respect to the relevant measure (typically $\mu$).
Theorem 3.9 ([2, Theorem 5.2]).
@@
The authors of [2] supply additional guarantees in Cases (2) and (3) of Theorem 3.9 (see [2, Section 4.3 and Corollaries A.3, A.4, A.5]). We summarize these guarantees in the following theorem.
Theorem 3.10 (Additional Guarantees for Theorem 3.9).
The following are additional guarantees for Cases (2) and (3) of Theorem 3.9. First, in both cases $\mu(A)\geq\frac{1}{4}\mu(V)$.
In Case (2), there exist constant $c>0$ and a sequence of cuts $S_{t}\subseteq A_{t}$, $0\leq t<T$ with $T=O(\log^{2}n)$, such that $A_{0}=V$, $A_{t+1}=A_{t}\setminus S_{t}$ (for $t\in\{0,\ldots T-1\}$). The returned cut is simply $(A,\bar{A})=(A_{T},\bigcup_{t=0}^{T-1}{S_{t}})$. The cuts apart from the last one satisfy $\mu\!\left(\bigcup_{t=0}^{T-2}{S_{t}}\right)=O\!\left(\frac{\mu(V)}{\log n}\right)$. Each $S_{t}$ is sparse: $\Phi^{\mu}_{G[A_{t}]}(A_{t+1},S_{t})\leq c\phi\log n$. Moreover, for every $t\in\{0,\ldots,T-1\}$,
  • There exists a flow routable with congestion $O(1)$ in $G[S_{t}]$, from the cut $(A_{t}\setminus S_{t},S_{t})$ to $S_{t}$, such that each vertex $v\in S_{t}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
  • There exists a flow routable with congestion $O(1)$ in $G[A_{t}\setminus S_{t}]$, from the cut $(A_{t}\setminus S_{t},S_{t})$ to $A_{t}\setminus S_{t}$, such that each vertex $v\in A_{t}\setminus S_{t}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
In other words, in each step we can route flow from the cut edges $E(A_{t}\setminus S_{t},S_{t})$ to either the side $S_{t}$ or $A_{t}\setminus S_{t}$.
In Case (3), this sequence of cuts exists as well, but a “trimming step” is run to get some larger cut $\bar{A}\supseteq\bigcup_{t=1}^{T-1}S_{t}$ which is returned. Moreover,
  • There exists a flow routable with congestion $O(1)$ in $G[A]$, from the cut $(A,\bar{A})$ to $A$, such that each vertex $v\in A$ receives at most $O(\phi\cdot\mu(v))$ units.
  • There exists a flow routable with congestion $O(1)$ in $G[\bar{A}]$, from the cut $(A,\bar{A})$ to $\bar{A}$, such that each vertex $v\in\bar{A}$ receives at most $O(\phi\log n\cdot\mu(v))$ units.
@@
In Case (3) of Theorem 3.9, the fact that $G[A]$ is an $\Omega(\phi)$-expander with respect to $\mu$, implies that AA $\Omega(\phi)$-expands with respect to $\mu$ in $G$ by Fact 3.7.
@@ 3.4 Demand States
When we describe the solution to a multi-commodity flow problem, it is useful to talk about intermediate stages of the flow. Analogously, we wish to talk about intermediate states of the distribution of demands when showing that a graph respects a given demand matrix. Intuitively, we imagine the vertices of the graph holding some demand of each commodity, which can be positive or negative. These demands move from one vertex to another according to demand matrices (i.e., multi-commodity flow problems), incurring some congestion (or, in the cut-sparsifier case, charge, as detailed in Section 5) on the edges of the graph. When a positive demand meets a negative demand of the same commodity, they cancel out. This scheme is made formal in the definitions in this section.
@@
Consider a set of commodities $\mathcal{K}$. The demand state of $v$, $v\in V$, is a vector, $P(v)$, where $P(v,k)$, $k\in\mathcal{K}$, defines the amount of commodity $k\in\mathcal{K}$ which resides at $v\in V$ and needs to be shipped out. Note that $P(v,k)$ may be either positive or negative, and in case it is negative, the commodity needs to be shipped into $v$. We denote by $\left\|P(v)\right\|_{1}=\sum_{k\in\mathcal{K}}{|P(v,k)|}$ the load on $v$.
A demand state $P$ of $G$ (or simply a demand state when $G$ is clear from context) is a collection of demand states, one for each $v\in V$, i.e., $P\in\mathop{\mathbb{R}}^{V\times\mathcal{K}}$, where the row of vertex $v$ is the demand state of $v$.
@@
$\operatorname{dem}_{P}(A,B)\mathrel{\mathop{:}}=\sum_{k\in\mathcal{K}}{\left|\sum_{u\in A}{P(u,k)}\right|}=\sum_{k\in\mathcal{K}}{\left|\sum_{u\in B}{P(u,k)}\right|}.$
@@
Throughout the paper we manipulate many demand states that are not themselves valid. Such demand states typically arise by restricting a valid demand state to a subgraph (or cluster), that is, by retaining only the demand vectors associated with vertices in the cluster. Although these restricted demand states may violate the validity condition, this poses no issue: they should be viewed as a partial or intermediate collection of demand vectors.
@@
Note that the notion of a valid demand state generalizes the notion of a demand matrix from Definition 2.1. Indeed, a demand matrix is a $V\times V$ matrix $Q$ where $Q(u,v)$ is the amount of $u$’s commodity that has to go to $v$. Such a specification corresponds to a valid demand state $P$ in which $\mathcal{K}=V$, the total amount shipped out of $u$ of its own commodity is $P(u,u)=\sum_{v\in V\setminus\{u\}}{Q(u,v)}$, and the amount shipped to $v$ of $u$’s commodity is P(v,u)=Q(u,v)P(v,u)=-Q(u,v), $v\neq u$ (note that the commodity is shipped into $v$, so $P(v,u)$ is negative). The standard definitions of congestion and demand across a cut of $Q$ equal those of $P$.
On the other hand, given a valid demand state, we can define the corresponding demand matrix (according to Definition 2.1) by adding to the graph a source vertex and sink vertex corresponding to each commodity, with appropriate additional edges capacitated according to the demands in $P$. The latter reduction is not be used in this paper, so we omit the full details.
@@
The following fact, which we use extensively, is a direct consequence of Definitions 3.13 and 3.15:
Fact 3.16.
@@
Suppose we have a demand state $P$. Given a demand matrix $Q$ (recall Definition 2.1), we can “move” the demand vectors in $P$ according to the requirements of $Q$. Each vertex $u\in V$ will send $Q(u,v)$ units in total to $v\in V$, and we split the demands evenly according to the distribution of commodities in the demand vector $P(u)$. In other words, $\frac{P(u,k)}{\left\|P(u)\right\|_{1}}\cdot Q(u,v)$ units of commodity $k\in\mathcal{K}$ will be shipped to $v$. If $P(u,k)<0$ this corresponds to shipping units of commodity $k$ from $v$ to $u$. This results in a demand state $P^{\uparrow Q}$, called the updated demand state:
@@
Example 3.18.
Suppose there are three commodities and the demand states of three vertices $u,v,w$ are
$P(u)=(20,-70,10),\qquad P(v)=(6,2,-2),\qquad P(w)=(-4,12,-4).$
Consider a demand matrix $Q$ that sends $30$ units from $u$ to $v$ and $10$ units from $w$ to $u$. Sending $30$ units from $u$ to $v$ moves $30\cdot(0.2,-0.7,0.1)=(6,-21,3)$ units of demand from $u$ to $v$. Sending $10$ units from $w$ to $u$ moves $10\cdot(-0.2,0.6,-0.2)=(-2,6,-2)$ units of demand from $w$ to $u$.
Applying both movements simultaneously, the updated demand vectors are
$P^{\uparrow Q}(u)=(12,-43,5),\qquad P^{\uparrow Q}(v)=(12,-19,1),\qquad P^{\uparrow Q}(w)=(-2,6,-2).$
We make frequent use of the following fact. It states that because we only move the mass around, no mass of any commodity is created or destroyed. This means that the total mass of each commodity is the same in $P$ and $P^{\uparrow Q}$, so it is zero in $P-P^{\uparrow Q}$. Therefore, $P-P^{\uparrow Q}$ is a valid demand state. The full proof of the fact is deferred to Appendix A.
@@
The next lemma presents two basic types of demand-state updates that will be used extensively throughout the analysis. The first corresponds to uniformly redistributing mass within a set, and the second to pushing all mass out of a set. Both updates arise frequently when we manipulate demand states.
Lemma 3.20.
Let $P$ be a (not necessarily valid) demand state on the set of commodities $\mathcal{K}$. Then:
@@
    Suppose $Q$ is a scaled all-to-all flow problem on some subset $A\subseteq V$, such that $Q(u,v)=\frac{\left\|P(u)\right\|_{1}}{|A|}$ if $u,v\in A$, and $Q(u,v)=0$ otherwise. Then, $P^{\uparrow Q}(v)=\frac{1}{|A|}\sum_{u\in A}{P(u)}$ for all $v\in A$, and $P^{\uparrow Q}(v)=P(v)$ for $v\notin A$. That is, $P^{\uparrow Q}$ distributes the mass in $A$ uniformly.
@@
    Suppose $Q$ moves demands from a set $A\subseteq V$ to $V\setminus A$: for all $u\in V$ and $v\in A$, $Q(u,v)=0$, and moreover, for all $u\in A$, $\sum_{v\in V\setminus A}{Q(u,v)}=\left\|P(u)\right\|_{1}$. In this case, $P^{\uparrow Q}$ is supported on $V\setminus A$.
Proof.
  1. 1.
    Let $v\in V$. If $v\notin A$, then $Q(v,u)=Q(u,v)=0$ for all $u\in V$, and therefore $P^{\uparrow Q}(v)=P(v)$. Assume now that $v\in A$, then
    $P^{\uparrow Q}(v)$ $=P(v)-\frac{P(v)}{\left\|P(v)\right\|_{1}}\cdot\sum_{u\in V}{Q(v,u)}+\sum_{u\in V}{\frac{P(u)}{\left\|P(u)\right\|_{1}}\cdot Q(u,v)}$
    $=P(v)-\frac{P(v)}{\left\|P(v)\right\|_{1}}\cdot\sum_{u\in A}{\frac{\left\|P(v)\right\|_{1}}{|A|}}+\sum_{u\in A}{\frac{P(u)}{\left\|P(u)\right\|_{1}}\cdot\frac{\left\|P(u)\right\|_{1}}{|A|}}$
    $=P(v)-P(v)+\frac{1}{|A|}\sum_{u\in A}{P(u)}$
    $=\frac{1}{|A|}\sum_{u\in A}{P(u)}.$
  2. 2.
    Let $v\in A$, then
    $P^{\uparrow Q}(v)$ $=P(v)-\frac{P(v)}{\left\|P(v)\right\|_{1}}\cdot\sum_{u\in V}{Q(v,u)}+\sum_{u\in V}{\frac{P(u)}{\left\|P(u)\right\|_{1}}\cdot Q(u,v)}$
    $=P(v)-\frac{P(v)}{\left\|P(v)\right\|_{1}}\cdot\sum_{u\in V\setminus A}{Q(v,u)}=0.$